OpenStudy (anonymous):

consider the question: x(t) = cos(2Pi*t) + sin(5t) is this system periodic? My work is below

OpenStudy (anonymous):

Where is your work?

OpenStudy (anonymous):

\[T = \frac{ 2\pi }{ \omega }\] \[T_1 = \frac{ 2\pi }{ 2\pi } = 1\] \[T_2 = \frac{ 2\pi }{ 5 }\] \[\frac{ T_1 }{ T_2 } = \frac{ 5 }{ 2\pi } \] since the last is irrational then its aperiodic?

OpenStudy (anonymous):

a system needs to satisfy x(t) = x(t+T) to be periodic since T1/T2 is irrational thus the system is aperiodic

OpenStudy (phi):

yes, that is correct.